OpenStudy (anonymous):

Evaluate each system below with the point (-1, -2/3). Select the system or systems which have the point (-1, -2/3) as a solution: a. -3x + 3y = 1 and y = -1 b. -2x + 6y = -2 and y = -2/3 c. x + y = -5/3 and x - y = -1/3 d. 3y = -2x and 2y = -3x e. -x + 3y = -2 and x = -1

OpenStudy (anonymous):

you need to plug the value of x or y from the point (-1,-2/3) into each equation. in choice c. , if you plug the value of x from the point, which is -1, the equation will be : (-1) + y = -5/3 -> y= -5/3 + 1 -> y= -2/3. and (-1) - y = -1/3 -> y = -1+ 1/3 -> y = -2/3 that means both equations from choice c have the point (-1, -2/3) therefore choice c is the right answer

OpenStudy (anonymous):

choice a, b and e can't be right because in choice a ,the equation y= -1 doesnt have any value of x in choice b, the equation y = -2/3 also doesnt have any value of x in choice e, the equation x =-1 doesnt have any value of y